OpenStudy (anonymous):

Suppose that g is a function with -x^2+6x-12  greater than or equal to g(x) less than or equal to x^2-6x+6 for all x. Evaluate lim x->3 g(x).

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} g(x)\] Suppose that g is a function with -x^2+6x-12 <or equal to g(x) <or equal to x^2-6x+6 for all x.

OpenStudy (anonymous):

let the first function be a and the 2nd be b then \[-x^2+6x-12 \le g(x) \le x^2-6x+6\] \[a \le g(x) \le b\] \[-3^2+6(3)-12 \le g(x) \le 3^2-6(3)+6\] \[-3 \le g(x) \le -3\] hence g(x) has to be 3 sandwich theorem